OpenStudy (anonymous):

need help solve the system of equations by graphing then classify the system as consistent or inconsistent, and as dependent or independent 4x-y=15 x+5y=-12 step by step plz

OpenStudy (anonymous):

First step, rewrite them as y=mx + b

OpenStudy (anonymous):

Check the consistency by matrices..

OpenStudy (anonymous):

These are linear equations in two variables If they have a common solution, then they will cross each other at a point and hence will be consistent. We put them in the form ax + by + c = 0 4x - y - 15 = 0 and x + 5y + 12 = 0 Now If \[a1/a2 \neq b1/b2\] then they will intersect at one point and will have a unique and consistent solution Here a1/a2 = 4/1 and b1/b2 = -1/5 Since a1/a2 is not equal to b1/b2, therefore it will have a unique solution and they are consistent.